Homework ditto that was handed out
Given f(x) = 3x + 1, find f(a + 3)

Answer by mathispowerful(115)


Given f(x) = 3x + 1, find f(a + 3)

here x = a+3, so we just replace x in 3x+1 by a+3
so we get f(a+3) = 3(a+3) + 1
= 3a+9+1
= 3a + 10